OpenStudy (younginslim):

Write the equation of the line that passes through (3, 4) and (2, –1) in slope-intercept form. y = 3x – 7 y = 3x – 5 y = 5x – 11 y = 5x – 9

OpenStudy (reemii):

With two points you can find the slope immediately, did you find it?

OpenStudy (younginslim):

3

OpenStudy (skullpatrol):

How did you calculate 3?

OpenStudy (younginslim):

idk

OpenStudy (skullpatrol):

So you guessed?

OpenStudy (younginslim):

I looked at the coordinates that was given

OpenStudy (reemii):

For the slope, it's "\(m = \frac{\Delta y}{\Delta x}\)", which means : (difference of y's) / (difference of x's).

OpenStudy (younginslim):

What that means

OpenStudy (younginslim):

The slope would y right?

OpenStudy (skullpatrol):

Nope

OpenStudy (skullpatrol):

Slope = rise/run Have you seen that^ formula before?

OpenStudy (younginslim):

Yes

OpenStudy (skullpatrol):

Good, call one point number 1 and the other number 2 and use $$\Huge slope=\dfrac{y_2-y_1}{x_2-x_1}$$

OpenStudy (younginslim):

I got it

OpenStudy (skullpatrol):

Good job :-D

OpenStudy (skullpatrol):

It's a good idea to make a sketch and plot the pionts as a check for your calculations.

OpenStudy (mathmale):

Now that you have the slope of this line, you need to write the actual equation of the line. Use the form y=mx+b. Here, m is the slope of the line and b is the y-intercept.